OpenStudy (anonymous):

1/4y – 3 = 9 (1 point)

OpenStudy (anonymous):

are u looking for the value of 'y'?.......if then 1/4y-3=9 cross 3 over and it becomes positive 1/4y=9+3=12 1/4y=12/1, cross multiply 1=4y*12(*=multiplication) 1=4*y*12=48*y 1=48*y, cross 48 over and it becomes division 1/48=y or y=1/48 or if the question is actually 1/(4y-3)=9 i.e if -3 is also being divided by 1 along with 4y, it would be 1/(4y-3)=9 cross over 4y-3 which becomes multiplied 1=9*(4y-3)=36y-27, cross -27 over 1+27=36y, 28=36y divide both sides by 36 28/36=36y/36 28/36=y, divide 28 and 36 by 4 y=7/9 depends on the question............................................................